# Phase Transition of Spectral Fluctuations in Large Gram Matrices with a Variance Profile: A Unified Framework for Sparse CLTs

### v1

@@ Abstract
We study the asymptotic spectral behavior of high-dimensional random Gram matrices with sparsity and a given variance profile, motivated by applications in wireless communication. Specifically, we consider the Gram matrices $\mathbf{S}_{n}=\mathbf{Y}_{n}\mathbf{Y}_{n}^{*}$, where the entries of $\mathbf{Y}_{n}$ are independent, centered, heteroscedastic, and sparse through Bernoulli masking. The sparsity level is parameterized as $s=q^{2}/n$, with $q$ ranging from polynomial order to order $n^{1/2}$.
We investigate two asymptotic regimes in a high-dimensional framework: a moderate-sparsity regime with fixed $s\in(0,1]$, and a high-sparsity regime where $s\to 0$. In both regimes, we establish the convergence of the empirical spectral distribution of $\mathbf{S}_{n}$ to a deterministic limit, and further derive central limit theorems for linear spectral statistics using resolvent techniques and martingale difference arguments. Our analysis reveals a phase transition in the fluctuation behavior across the two regimes. In the high-sparsity regime, the asymptotic fluctuations are governed by fourth-moment effects, with sparsity-scaled contributions being suppressed. Moreover, a mismatch between the scaling of the mean and variance, of different orders in $q$, necessitates an explicit correction in the centering of the linear spectral statistic. The theory applies to both Gaussian and non-Gaussian entries, and its statistical utility is illustrated through applications to hypothesis testing and outage probability analysis in large-scale MIMO systems.
@@ 1 Introduction
A primary motivation arises from wireless communications. In multiple-input multiple-output (MIMO) channels, $\mathbf{Y}_{n}$ may represent the channel matrix, where $b_{ij}$ models random link availability and $\sigma_{ij}^{2}$ characterizes non-uniform propagation effects. The associated Gram matrix $\mathbf{S}$ determines essential performance metrics, such as the mutual information $C(\sigma^{2})=\log\det(1+\mathbf{S}/\sigma^{2})$ with a given parameter $\sigma^{2}>0$, and its fluctuation governs the outage probability—the probability that $C(\sigma^{2})$ falls below a target transmission rate. The study of linear spectral statistics (LSS) of $\mathbf{S}$ provides a natural probabilistic framework for quantifying these fluctuations; see, for instance, [11, 9, 5, 12, 30, 28]. Consequently, central limit theorems (CLTs) for LSS of Gram-type matrices have become indispensable tools in the stochastic analysis of mutual-information variations.
Beyond capacity analysis, similar ideas facilitate statistical inference on large-scale fading structures across heterogeneous environments. In sparse MIMO systems, the channel gain can be modeled as $y_{ij}=b_{ij}l_{ij}g_{ij}/\sqrt{ns}$, where $\{l_{ij}\}$ represent large-scale fading coefficients, $\{g_{ij}\}$ are independent and identically distributed (i.i.d.) small-scale fading. The matrix $\mathbf{L}=(l_{ij})$ varies slowly with the macroscopic environment, and testing whether two scenarios share the same large-scale fading pattern offers a approach to detecting environmental or topological changes [21, 17, 22]. In this context, CLTs for LSS of sparse and heteroscedastic Gram matrices serve as a unified framework linking random matrix theory with practical inference tasks in high-dimensional sparse MIMO systems.
@@
    the high-sparsity regime, where $s=q^{2}/n\to 0$ with $q=n^{\phi}$ and $n^{\phi}\leq q\leq n^{1/2}$ for some fixed $\phi>0$.
These two regimes jointly characterize the transition from moderately sparse to extremely sparse structures. For both regimes, we first establish the convergence of the empirical spectral distribution of $\mathbf{S}_{n}$ to a nonrandom limiting spectral distribution (LSD). Building on this foundation, we then derive CLTs for the LSS using a martingale difference approach. The mean and variance terms are determined through the unique solutions of certain systems of linear functional equations. When proving the CLT, additional care is required: the analysis necessitates a fourth-order Lindeberg condition, which holds under the existence of moments of order $8+\epsilon$ for some small $\epsilon>0$.
The contributions of this paper can be summarized as follows. First, we identify a clear phase transition in the fluctuation behavior between the two sparsity regimes. Although the high-sparsity CLT can be viewed as a limiting case of the moderate-sparsity regime as $s\to 0$, the fluctuation structure changes qualitatively: certain terms vanish, and the limiting distribution is governed solely by the remaining ones. In particular, in the high-sparsity regime, the CLT is dominated by terms associated with the fourth-order moments of the entries, while the fluctuation components scaled by the sparsity level are suppressed. Second, we uncover a mismatch in convergence rates for the mean and variance terms of the centered LSS. The variance scales as $\sqrt{p}q$, while the mean scales as $q^{2}$. This mismatch is asymptotically negligible in the moderate-sparsity regime but causes divergence of the mean in the high-sparsity case. To address this issue, we introduce a centering correction by subtracting the diverging mean term, thereby obtaining a valid limiting distribution. Third, we develop a unified and general framework applicable to both Gaussian and non-Gaussian entries, and extending beyond the logarithmic function to a broad class of analytic test functions.
@@ 2.1 The model
Let $p=p(n)$ be a sequence of integers such that $\lim_{n\rightarrow\infty}{p}/{n}=c\in(0,\infty)$. Consider a $p\times n$ random matrix $\mathbf{Y}_{n}=(y_{ij})$ with entries
@@
Assume that $(w_{ij};1\leq i\leq p,1\leq j\leq n)$ are real or complex independent distributed (i.i.d.) random variables, satisfying the moment conditions
@@
The lower bound in Assumption 2.3 ensures mild non-degeneracy of the column-wise average variances. Although a nontrivial limiting spectral distribution can be obtained under weaker averaged conditions [10], this assumption is crucial for stable resolvent estimates and is commonly imposed in central limit theorems [11].
@@
Assume that $(b_{ij};1\leq i\leq p,1\leq j\leq n)$ are i.i.d. Bernoulli random variables with parameter $s$, independent of $(w_{ij})$. The sparsity parameter $q$ is introduced through $s=q^{2}/n$ with $q=n^{\phi}$, where $\phi>0$.
Define the matrix $\mathbf{X}_{n}\in\mathbb{R}^{p\times n}$ with the entries $x_{ij}={b_{ij}w_{ij}}/{\sqrt{s}}$. Writing $\mathbf{Y}_{n}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})$ and $\mathbf{X}_{n}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})$ with $\mathbf{y}_{j}=\mathbf{x}_{j}/{\sqrt{n}}\in\mathbb{R}^{p}$, the Gram matrix can be expressed as
@@
where $\circ$ denotes the Hadamard (elementwise) product, $\mathbb{B}=(b_{ij})$, and $\mathbf{W}_{n}=(w_{ij})$. Equivalently,
@@ 2.2 Notations
For matrices, $\operatorname{diag}(a_{i};1\leq i\leq k)$ denotes the $k\times k$ diagonal matrix with entries $a_{1},\dots,a_{k}$. For a matrix $\mathbf{A}$, its $(i,j)$-th entry is denoted by either $a_{ij}$ or $\mathbf{A}]_{ij}$, depending on context. We write $\mathbf{A}^{T}$ for the transpose, $\mathbf{A}^{*}$ for the conjugate transpose, $\operatorname{Tr}(\mathbf{A})$ for the trace, and $\det(\mathbf{A})$ for the determinant when $\mathbf{A}$ is square. The empirical spectral distribution (ESD) of $\mathbf{A}\mathbf{A}^{*}$ is defined by
@@ 3 Main results
3.1 The limiting spectra distribution
@@
$m(z):=\int_{\mathbb{R}}\frac{\mathrm{d}\nu(x)}{x-z}\quad\left(z\in\mathbb{C}\setminus\mathbb{R}^{+}\right).$
@@
The resolvent of a matrix plays a central role in random matrix theory, as it is closely linked to the Stieltjes transform of ESD. For any $z\in\mathbb{C}\setminus\mathbb{R}^{+}$, we define
@@
Let $(\mathbf{Y}_{n}\mathbf{Y}_{n}^{*})$ be a family of random matrices satisfying Assumptions 2.12.2 and 2.4, where Assumption 2.1 can be relaxed to require only a finite moment of order $4+\epsilon$ for some $\epsilon>0$. Define the system of $p+n$ equations:
$\begin{cases}t_{i}(z)=\dfrac{-1}{z\left(1+(1/n)\operatorname{Tr}\left(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}_{n}(z)\right)\right)},&\text{for }1\leq i\leq p,\\[10.0pt] \tilde{t}_{j}(z)=\dfrac{-1}{z\left(1+(1/n)\operatorname{Tr}\left(\mathbf{\Sigma}_{j}\mathbf{T}_{n}(z)\right)\right)},&\text{for }1\leq j\leq n,\end{cases}$ (3)
@@ 3.2 CLT in the moderate-sparsity regime
This section focuses on the moderate-sparsity regime, where the retention probability $s\in(0,1]$ remains constant as $n\to\infty$. To study the second-order fluctuations of the LSS, we introduce the centered and normalized statistic
$L^{c}(f,\mathbf{S}_{n})=\sqrt{p}\,q\int f(x)\,\mathrm{d}\!\left[F^{\mathbf{S}_{n}}(x)-\pi_{n}(x)\right],$
where $\pi_{n}$ denotes the deterministic equivalent of the ESD $F^{\mathbf{S}_{n}}$. The scaling factor $\sqrt{p}q$ reflects the variance structure under sparsity. While the classical $n$-scaling remains applicable when $s$ is fixed (since $\sqrt{p}q$ is of the same order as $n$), we therefore employ $\sqrt{p}q$ for a unified treatment across both moderate- and high-sparsity regimes. This choice is necessary because, as shown in equation (1) of the supplementary materials, it is the only scaling that yields a non-degenerate limit in the high-sparsity case.
@@
Let $f_{1},\dots,f_{k}$ be functions that are analytic in a complex neighborhood of the support of the LSD of $\mathbf{S}_{n}$, and continuous at zero. Suppose Assumptions 2.12.4 hold. Then the random vector $\left(L^{c}(f_{j},\mathbf{S}_{n})\right)_{j=1}^{k}$ is tight and converges weakly to a Gaussian vector $\left(X_{f_{j}}\right)_{j=1}^{k}$. The mean and covariance structure of the limit are given as follows.
@@
    where $\mathscr{E}_{n}(z)=\frac{1}{n}\sum_{j=1}^{n}\psi_{j,n}(z)$, and $\{\psi_{j,n}(z)\}_{j=1}^{n}$ is the unique solution to the system
    $\psi_{j}(z)=\sum_{m=1}^{n}a_{jm}\psi_{m}+\theta_{j}(z),\quad 1\leq j\leq n.$
@@
    $+\frac{q}{\sqrt{p}}\frac{1}{n}z^{3}\tilde{t}_{j}^{2}(z)\sum_{i=1}^{p}\sigma_{ij}^{2}t_{i}^{3}(z)\left\{\kappa\tilde{U}_{i}-\frac{\kappa+2}{n}\operatorname{Tr}\left(\widetilde{\mathbf{T}}(z)\circ\widetilde{\mathbf{T}}(z)\circ\widetilde{\mathbf{\Sigma}}_{i}^{2}\right)\right\}$
    $+\frac{q}{\sqrt{p}}z^{2}\tilde{t}_{j}^{3}(z)\left\{\kappa U_{j}-\frac{\kappa+2}{n}\operatorname{Tr}\left(\mathbf{T}(z)\circ\mathbf{T}(z)\circ\mathbf{\Sigma}_{j}^{2}\right)\right\},$
    with $(\tilde{U}_{i})_{1\leq i\leq p}$ and $(U_{j})_{1\leq j\leq n}$ determined by the systems in Lemma A.1.
@@
    $\mathscr{V}_{n}(z_{1},z_{2})=\frac{\partial^{2}}{\partial z_{2}\partial z_{1}}$ $\left\{\frac{s(\kappa+1)}{p}\sum_{j=1}^{n}H_{jj}\right$.
@@
    Here, $\mathbf{T}^{(1)}(z)=\left(\mathbf{I}+\frac{1}{n}\sum_{k=1}^{n}\tilde{t}_{k}(z)\mathbf{\Sigma}_{k}\right)^{-1}$, and for each $j$, $\{H_{\ell j}\}_{\ell=1}^{j}$ is the unique solution to the triangular system
    $H_{\ell j}=\sum_{i=1}^{j-1}a_{\ell i}H_{ij}+na_{\ell j},\quad 1\leq\ell\leq j.$
@@
To verify the consistency of our CLT with those established for complete data in earlier studies, consider the special case $\mathbf{\Sigma}=\sigma^{2}\mathbf{1}_{p\times p}$ and $s=1$. Then, the model in (2) becomes $\mathbf{y}_{j}=\mathbf{\Gamma}\mathbf{z}_{j}/\sqrt{n}$, where $\mathbf{\Gamma}=\sigma^{2}\mathbf{I}_{p}$, and $(z_{ij})$ are i.i.d. random variables with mean zero and unit variance. The LSD of $\mathbf{\Gamma}$ is the point mass $H(t)=\delta_{\sigma^{2}}(t)$. In this case, $\tilde{t}_{j}(z)$ is identical for all $j$ and is denoted by $\underline{m}(z)$, which satisfies

### v2

@@ Abstract
We study the asymptotic spectral behavior of high-dimensional random Gram matrices with sparsity and a variance profile, motivated by applications in wireless communications. Specifically, we consider the Gram matrices $\mathbf{S}_{n}=\mathbf{Y}_{n}\mathbf{Y}_{n}^{*}$, where the entries of $\mathbf{Y}_{n}$ are independent, centered, heteroscedastic, and sparse through Bernoulli masking. The sparsity level is parameterized as $s=q^{2}/n$, where $q$ ranges from polynomial order up to order $n^{1/2}$.
We investigate two asymptotic regimes: a moderate-sparsity regime with fixed $s\in(0,1]$, and a high-sparsity regime where $s\to 0$. In both regimes, we establish the convergence of the empirical spectral distribution of $\mathbf{S}_{n}$ to a deterministic limit, and further derive central limit theorems for linear spectral statistics using resolvent techniques and martingale difference arguments. Our analysis reveals a phase transition in the fluctuation behavior across the two regimes. In the high-sparsity regime, the asymptotic fluctuations are entirely governed by fourth-moment effects, with sparsity-scaled contributions being suppressed. Moreover, the leading deterministic term and the variance of the linear spectral statistic scale at different rates in $q$, causing the standard centering to fail and necessitating an explicit correction to recover a valid CLT. The results apply to both Gaussian and non-Gaussian entries and are illustrated through applications to hypothesis testing and outage probability analysis in large-scale MIMO systems.
@@ 1 Introduction
A primary motivation arises from wireless communications. In multiple-input multiple-output (MIMO) channels, $\mathbf{Y}_{n}$ may represent the channel matrix, where $b_{ij}$ models random link availability and $\sigma_{ij}^{2}$ characterizes non-uniform propagation effects. The associated Gram matrix $\mathbf{S}$ determines essential performance metrics, such as the mutual information $C(\sigma^{2})=\log\det(1+\mathbf{S}/\sigma^{2})$ with a given parameter $\sigma^{2}>0$, and its fluctuation governs the outage probability—the probability that $C(\sigma^{2})$ falls below a target transmission rate. The study of linear spectral statistics (LSS) for $\mathbf{S}$ provides a natural probabilistic framework for quantifying these fluctuations; see, for instance, [11, 9, 5, 12, 29, 30]. Consequently, central limit theorems (CLTs) for LSS of Gram-type matrices have become indispensable tools in the stochastic analysis of mutual-information variations.
Beyond capacity analysis, similar ideas facilitate statistical inference on large-scale fading structures across heterogeneous environments. In sparse MIMO systems, the channel gain can be modeled as $y_{ij}=b_{ij}l_{ij}g_{ij}/\sqrt{ns}$, where $\{l_{ij}\}$ represent large-scale fading coefficients, $\{g_{ij}\}$ are independent and identically distributed (i.i.d.) small-scale fading. The matrix $\mathbf{L}=(l_{ij})$ varies slowly with the macroscopic environment, and testing whether two scenarios share the same large-scale fading pattern offers an approach to detecting environmental or topological changes [21, 17, 22]. In this context, CLTs for LSS of sparse and heteroscedastic Gram matrices serve as a unified framework linking random matrix theory with practical inference tasks in high-dimensional sparse MIMO systems.
@@
    the high-sparsity regime, where $s=q^{2}/n\to 0$ with $n^{\phi}\leq q<n^{1/2}$ for some fixed $\phi>0$.
These two regimes jointly characterize the transition from moderately sparse to extremely sparse structures. For both regimes, we first establish the convergence of the empirical spectral distribution (ESD) of $\mathbf{S}_{n}$ to a nonrandom limiting spectral distribution (LSD). Building on this foundation, we then derive CLTs for the LSS using a martingale difference approach. The mean and variance terms are determined through the unique solutions of certain systems of linear functional equations. When proving the CLT, additional care is required: the analysis necessitates a fourth-order Lindeberg condition, which holds under the existence of moments of order $8+\epsilon$ for some small $\epsilon>0$.
The contributions of this paper can be summarized as follows. First, we identify a clear phase transition in the fluctuation behavior between the two sparsity regimes. While the high-sparsity CLT can be viewed as a limiting case of the moderate-sparsity regime as $s\to 0$, the fluctuation structure changes qualitatively: certain components vanish, and the limiting distribution is determined solely by the remaining terms. In particular, in the high-sparsity regime, the limiting Gaussian fluctuations are dominantly driven by fourth-order moment effects, while contributions arising from sparsity-dependent terms vanish asymptotically. Second, we uncover a mismatch in convergence rates for the mean and variance terms of the centered LSS. After normalization by $\sqrt{p}q$, the variance remains of order one, whereas the corresponding mean term grows at rate $\sqrt{p}/q$, diverging in the high-sparsity regime. While this effect is asymptotically negligible under moderate sparsity, it invalidates the CLT when $q$ is too small. To address this issue, we introduce a centering correction by subtracting the diverging mean term, thereby obtaining a valid limiting distribution. Third, we develop a unified and general framework applicable to both Gaussian and non-Gaussian entries, and extending beyond the logarithmic function to a broad class of analytic test functions.
@@ 2.1 The model
Let $p=p(n)$ be a sequence of integers such that $\lim_{n\rightarrow\infty}{p}/{n}=c\in(0,\infty)$. Consider a $p\times n$ real- or complex-valued random matrix $\mathbf{Y}_{n}=(y_{ij})$ with entries
@@
Assume that $(w_{ij};1\leq i\leq p,1\leq j\leq n)$ are real- or complex-valued independently distributed random variables, satisfying the moment conditions
@@
The lower bound in Assumption 2.3 ensures mild non-degeneracy of the column-wise average variances. Although a nontrivial LSD can be obtained under weaker conditions [10], this assumption is crucial for stable resolvent estimates and is commonly imposed in CLTs [11].
@@
Assume that $(b_{ij};1\leq i\leq p,1\leq j\leq n)$ are i.i.d. Bernoulli random variables with parameter $s$, independent of $(w_{ij})$. The sparsity parameter $q$ is introduced through $s=q^{2}/n$ with $n^{\phi}\leq q\leq n^{1/2}$ for some fixed $\phi>0$.
Define the matrix $\mathbf{X}_{n}$ with the entries $x_{ij}={b_{ij}w_{ij}}/{\sqrt{s}}$. Writing $\mathbf{Y}_{n}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})$ and $\mathbf{X}_{n}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})$ with $\mathbf{y}_{j}=\mathbf{x}_{j}/{\sqrt{n}}$, the Gram matrix can be expressed as
@@
where $\circ$ denotes the Hadamard (element-wise) product, $\mathbb{B}=(b_{ij})$, and $\mathbf{W}_{n}=(w_{ij})$. Equivalently,
@@ 2.2 Notations
For matrices, $\operatorname{diag}(a_{i};1\leq i\leq k)$ denotes the $k\times k$ diagonal matrix with entries $a_{1},\dots,a_{k}$. For a matrix $\mathbf{A}$, its $(i,j)$-th entry is denoted by either $a_{ij}$ or $\mathbf{A}]_{ij}$, depending on context. We write $\mathbf{A}^{T}$ for the transpose, $\mathbf{A}^{*}$ for the conjugate transpose, $\operatorname{Tr}(\mathbf{A})$ for the trace, and $\det(\mathbf{A})$ for the determinant when $\mathbf{A}$ is square. The ESD of $\mathbf{A}\mathbf{A}^{*}$ is defined by
@@ 3 Main results
3.1 The limiting spectral distribution
@@
$m_{v}(z):=\int_{\mathbb{R}}\frac{\mathrm{d}v(x)}{x-z}\quad\left(z\in\mathbb{C}\setminus\mathbb{R}^{+}\right).$
@@
The resolvent of a matrix plays a central role in random matrix theory, as it is closely linked to the Stieltjes transform of the ESD. For any $z\in\mathbb{C}\setminus\mathbb{R}^{+}$, we define
@@
Let $\{\mathbf{Y}_{n}\mathbf{Y}_{n}^{*}\}$ be a family of random matrices satisfying Assumptions 2.12.2 and 2.4. Moreover, Assumption 2.1 can be relaxed to require only that $w_{ij}$ have uniformly bounded $(4+\epsilon)$-th moments for some $\epsilon>0$. Define the system of $p+n$ equations:
$\begin{cases}t_{i}(z)=\dfrac{-1}{z[1+(1/n)\operatorname{Tr}(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}_{n}(z))]},&\text{for }1\leq i\leq p,\\ \tilde{t}_{j}(z)=\dfrac{-1}{z\left[1+(1/n)\operatorname{Tr}\left(\mathbf{\Sigma}_{j}\mathbf{T}_{n}(z)\right)\right]},&\text{for }1\leq j\leq n,\end{cases}$ (3)
@@ 3.2 CLT in the moderate-sparsity regime
This section focuses on the moderate-sparsity regime, where the retention probability $s\in(0,1]$ remains constant as $n\to\infty$. To capture non-degenerate second-order fluctuations, we consider the centered and properly normalized version of the LSS,
$L^{c}(f,\mathbf{S}_{n})=\sqrt{p}\,q\int f(x)\,\mathrm{d}\!\left[F^{\mathbf{S}_{n}}(x)-\pi_{n}(x)\right],$ (4)
where $\pi_{n}$ denotes the deterministic equivalent of the ESD $F^{\mathbf{S}_{n}}$. The scaling factor $\sqrt{p}q$ reflects the variance structure induced by sparsity. While the classical $n$-scaling remains applicable when $s$ is fixed (since $\sqrt{p}q$ is of the same order as $n$), we therefore employ $\sqrt{p}q$ for a unified treatment across both moderate- and high-sparsity regimes. This choice is necessary because, as shown in equation (1) of the supplementary material, it is the only scaling that yields a non-degenerate limit in the high-sparsity case.
@@
Let $f_{1},\dots,f_{k}$ be functions that are analytic in a complex neighborhood of the support of the LSD of $\mathbf{S}_{n}$, and continuous at zero. Suppose Assumptions 2.12.4 hold and $s\in(0,1]$ is fixed. Then the random vector $\left(L^{c}(f_{j},\mathbf{S}_{n})\right)_{j=1}^{k}$ is tight and converges weakly to a Gaussian vector $\left(X_{f_{j}}\right)_{j=1}^{k}$. The mean and covariance structure of the limit are given as follows.
@@
    where $\mathscr{E}_{n}(z)=\frac{1}{n}\sum_{j=1}^{n}\psi_{j}(z)$, and $\{\psi_{j}(z)\}_{j=1}^{n}$ is the unique solution to the system
    $\psi_{j}(z)=\sum_{m=1}^{n}a_{jm}(z,z)\psi_{m}(z)+\theta_{j}(z),\quad 1\leq j\leq n.$
@@
    $+\frac{q}{\sqrt{p}}\frac{1}{n}z^{3}\tilde{t}_{j}^{2}(z)\sum_{i=1}^{p}\sigma_{ij}^{2}t_{i}^{3}(z)\left\{\kappa\tilde{U}_{i}(z)-\frac{\kappa+2}{n}\operatorname{Tr}\left(\widetilde{\mathbf{T}}(z)\circ\widetilde{\mathbf{T}}(z)\circ\widetilde{\mathbf{\Sigma}}_{i}^{2}\right)\right\}$
    $+\frac{q}{\sqrt{p}}z^{2}\tilde{t}_{j}^{3}(z)\left\{\kappa U_{j}(z)-\frac{\kappa+2}{n}\operatorname{Tr}\left(\mathbf{T}(z)\circ\mathbf{T}(z)\circ\mathbf{\Sigma}_{j}^{2}\right)\right\},$
    with $(\tilde{U}_{i}(z))_{1\leq i\leq p}$ and $(U_{j}(z))_{1\leq j\leq n}$ determined by the systems in Lemma A.1.
@@
    $\mathscr{V}_{n}(z_{1},z_{2})=\frac{\partial^{2}}{\partial z_{2}\partial z_{1}}$ $\left\{\frac{s(\kappa+1)}{p}\sum_{j=1}^{n}H_{jj}(z_{1},z_{2})\right$.
@@
    Here, $\mathbf{T}^{(1)}(z)=\left(\mathbf{I}+\frac{1}{n}\sum_{k=1}^{n}\tilde{t}_{k}(z)\mathbf{\Sigma}_{k}\right)^{-1}$, and for each $j$, $\{H_{\ell j}(z_{1},z_{2})\}_{\ell=1}^{j}$ is the unique solution to the triangular system
    $H_{\ell j}(z_{1},z_{2})=\sum_{i=1}^{j-1}a_{\ell i}(z_{1},z_{2})H_{ij}(z_{1},z_{2})+na_{\ell j}(z_{1},z_{2}),\quad 1\leq\ell\leq j.$
@@
To verify the consistency of our CLT with those established for complete data in earlier studies, consider the special case where the variance profile matrix $\mathbf{\Sigma}:=(\sigma_{ij}^{2})$ reduces to $\mathbf{\Sigma}=\sigma^{2}\mathbf{1}_{p\times p}$ and $s=1$. Then, the model in (2) becomes $\mathbf{y}_{j}=\mathbf{\Gamma}\mathbf{z}_{j}/\sqrt{n}$, where $\mathbf{\Gamma}=\sigma^{2}\mathbf{I}_{p}$, and $(z_{ij})$ are i.i.d. random variables with mean zero and unit variance. The LSD of $\mathbf{\Gamma}$ is the point mass $H(t)=\delta_{\sigma^{2}}(t)$. In this case, $\tilde{t}_{j}(z)$ is identical for all $j$ and is denoted by $\underline{m}(z)$, which satisfies